OpenStudy (anonymous):

what is y if y=x+3times 8\12?

OpenStudy (anonymous):

hi feather

OpenStudy (z4k4r1y4):

lol

OpenStudy (z4k4r1y4):

i prefer quill

OpenStudy (anonymous):

what your picture is a picture of a feather(quill)

OpenStudy (anonymous):

quill,what is your real name?

OpenStudy (anonymous):

lol

OpenStudy (z4k4r1y4):

we have two unknowns. we need to know what x equals to find 'y'

OpenStudy (anonymous):

yes

OpenStudy (z4k4r1y4):

from what you've posted there isn't enough information.

OpenStudy (z4k4r1y4):

unless the answer is just y = x + 3*8/12 then the answer is y = x + 2